# Bayesian Differentiable Physics for Cloth Digitalization

### 27 Feb 2024 v1

@@ 3.2 A Bayesian Model for Cusick Drape
Figure 2: (a) A circular (diameter=$30cm$) cloth sample for Cusick drape test. (b) Cloth sample mesh generated in FreeFEM++[33] by Delaunay triangulation which has 2699 vertices, 7924 edges (7754 bending edges), and 5226 faces. A bending edge (highlighted in orange) is shared between two adjacent triangles, so the edges highlighted in blue (the boundary) are not bending edges.
We discretize a cloth sample into a triangular mesh with $v$ vertices (Fig. 2(b)). Then we define the sample’s state as $\mathcal{S}_{t}=\{\mathbf{x}_{t},\dot{\mathbf{x}}_{t}\}$ where $\mathbf{x}_{t}\in\mathbb{R}^{3v}$ and $\dot{\mathbf{x}}_{t}\in\mathbb{R}^{3v}$ are the vertex position and velocity respectively at time $t$. Therefore, a draping motion with discretized time is $\mathcal{S}_{0:n}=\{\mathcal{S}_{t}:t\in\mathbb{Z}^{+};t\leq n\}$, with a time step size $h$. Since we only observe the final image $\mathcal{I}$ and the initial state $\mathcal{S}_{0}$, their joint probability is:
$p(\mathcal{I},\mathcal{S}_{0})=$⋯∫$\idotsint p(\mathcal{I}|\mathcal{S}_{n},\tau)∫⋯$
$\prod_{i=0}^{n-1}p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)p(\tau)d\mathcal{S}_{1:n}d\tau$ (1)
@@
Sec. 3.2 is not easy to estimate due to two challenges. First, unlike the prior works which depend on dense observations on the intermediate state transitions [79] to estimate $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$, Cusick drape testing does not capture the cloth sample’s motion. Also, we do not observe the full $\mathcal{S}_{n}$, but only its (simplified) 2D representation $\mathcal{I}$.
@@
To this end, we assume two deterministic mappings can be established for $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$ and $p(\mathcal{I}|\mathcal{S}_{n},\tau)$. The determinism assumptions are reasonable as $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$ can be seen as a quasi-deterministic physical process, subject to minor system stochasticity which is largely mitigated by the rigorous control of Cusick settings. $p(\mathcal{I}|\mathcal{S}_{n},\tau)$ can be seen as a rendering process studied in computer graphics. Here we are mainly interested in the silhouette, following the practice in textile in Cusick drape analysis [16]. Therefore, we replace $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$ with a state transition function $\mathcal{S}_{i+1}=s(\mathcal{S}_{i},\tau)$ where $s$ is a deterministic function then $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)=p(s(\mathcal{S}_{i},\tau)|\mathcal{S}_{i},\tau)=1$. Similarly, $p(\mathcal{I}|\mathcal{S}_{n},\tau)=p(r(\mathcal{S}_{n})|\mathcal{S}_{n},\tau)=1$ where $r$ is a rendering function. Therefore, Sec. 3.2 is transformed to:
@@ 4 Data Collection
The Cusick drape data is collected following the BS EN ISO 9073-9:2008 [1]. In every test, the warp/wale and weft/course directions are aligned across samples to ensure the same initial condition. In addition, we are able to reconstruct the 3D meshes of cloth drapes thanks to our patented drape meter. However, the 3D data are mainly for evaluation and the silhouette images are for learning. This is because not every Cusick drape meter can reconstruct 3D meshes out of cloth drapes. Being able to learn from the 2D silhouette images only is crucial in making our method applicable in real-world settings. In addition, we also measure the sample weight and calculate the average area density $\rho=(\sum_{i<13}^{i=1}m_{i})/(12\pi R^{2})$ where $m_{i}$ denotes the measured weight of cloth sample $i$ and $R=0.15m$. We also measure the thickness and include them in our dataset. More details are in the SM.
@@ 5 Experiments
Due to the nature of limited data, we conduct our experiments on small datasets. We use 5 representative types of cloths (each with 12 samples) that exhibit visually distinguishable drapability. For convenience, we name cloths as “material color”, e.g. Cotton Blue, Cotton Pink. To digitalize a cloth, we randomly select 1 out of 12 samples for training to learn their parameter distributions. During learning, we run 100 steps ($n=100$) for the forward simulation, with time step size $h=0.05s$ and the total time lapse is $5s$. To qualitatively evaluate the digitalization, we simulate garments made from different digitalized cloths and visually compare the simulated garments with their corresponding real cloths. To evaluate the model (i.e. drape fitting), we employ several metrics. For fitting capability, we use mean squared error (MSE) between the fitted and the ground truth drape images. We also use Hausdorff distance (H.Dis) between the simulated 3D mesh and the ground-truth (GT) mesh. Further, we use radius-angle graph (Fig. 6(a)[44] which is widely adopted for describing Cusick drape waves and comparing drape shapes. As generalization, we also test if our model learned on one sample can predict the physics of unseen cloth samples of the same type. Please refer to https://youtu.be/ProN0y1bURY for more visual results.
@@ 5.2 Comparison
To our best knowledge, there is no similar method designed for exactly the same setting as ours, we therefore adopt the closest methods as baselines. We adopt [51] because their method can also digitalize cloth by learning cloth physical parameters, albeit only taking simulation data as input. Since they only model homogeneous material, we refer to their model as HOMO. Further, we augment their model by making the cloth physical parameter learning element-wise, to enable learning heterogeneous materials. We call this model HETER for short. Both HOMO and HETER are deterministic models. So we use MSE as the loss function: $\mathcal{L}(\tau)=\sum_{i=1}^{L}\sum_{j=1}^{L}(\hat{\mathcal{I}}_{ij}-\mathcal{I}_{ij})^{2}$, for training. For comparison, we train HOMO and HETER with the same data (1 for training and the rest 11 for testing) as BDP. For BDP, we draw parameters from the learned parameter distributions for 1000 times and run Cusick drape simulations. Then, we select the best result to calculate the MSE and H. Dis.
@@
While our method is built on differentiable physics models for learning, i.e. derivative-based optimization, the traditional methods widely used in material science and physics are usually based on derivative-free optimization. So we also compare different learning strategies.
Bayesian Optimization (BO) is a representative derivative-free optimizer which usually uses the Gaussian Process to approximate an unknown optimized objective function [12, 25]. However, the performance of vanilla BO drops drastically when the number of parameters is above 20 [24, 46]. There are over 200,000 parameters in the HETER cloth model. So we employ Random Embedding Bayesian Optimization (REMBO) [76] and Hashing-enhanced Subspace Bayesian Optimization (HeSBO) [58] as baselines. We use BoTorch’s [2] REMBO and HeSBO implementation and compare them with our BDP. Given the same drape silhouette, we run our method 1000 epochs (gain the best optimization result within 200 epochs), and the REMOB and HeSBO for 500 trials. Tab. 2 shows that, our derivative-based method is better with fewer optimization steps. More results can be found in the SM.

### 08 Mar 2024 v3

@@ 3.2 A Bayesian Model for Cusick Drape
Figure 2: (a) A circular (diameter=$30cm$) cloth sample for Cusick drape test. (b) Cloth sample mesh has 2699 vertices, 7924 edges (7754 bending edges), and 5226 faces. A bending edge (highlighted in orange) is shared between two adjacent triangles, so the edges highlighted in blue (the boundary) are not bending edges.
We discretize a cloth sample into a triangular mesh with $v$ vertices (Fig. 2(b)). Then we define the sample’s state as $\mathcal{S}_{t}=\{\mathbf{x}_{t},\dot{\mathbf{x}}_{t}\}$ where $\mathbf{x}_{t}\in\mathbb{R}^{3v}$ and $\dot{\mathbf{x}}_{t}\in\mathbb{R}^{3v}$ are the vertex position and velocity respectively at time $t$. Therefore, a draping motion with discretized time is $\mathcal{S}_{0:n}=\{\mathcal{S}_{t}:t\in\mathbb{Z}^{+};t\leq n\}$, with a time step size $h$. Since we only observe the final image $\mathcal{I}$ and the initial state $\mathcal{S}_{0}$, their joint probability $p(\mathcal{I},\mathcal{S}_{0})=$
∫⋯∫$\idotsint p(\mathcal{I}|\mathcal{S}_{n},\tau)\prod_{i=0}^{n-1}p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)p(\tau)d\mathcal{S}_{1:n}d\tau∫⋯$ (1)
@@
Eq. 1 is not easy to estimate due to two challenges. First, unlike the prior works which depend on dense observations on the intermediate state transitions [78] to estimate $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$, Cusick drape testing does not capture the cloth sample’s motion. Also, we do not observe the full $\mathcal{S}_{n}$, but only its (simplified) 2D representation $\mathcal{I}$.
@@
To this end, we assume two deterministic mappings can be established for $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$ and $p(\mathcal{I}|\mathcal{S}_{n},\tau)$. The determinism assumptions are reasonable as $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$ can be seen as a quasi-deterministic physical process, subject to minor system stochasticity which is largely mitigated by the rigorous control of Cusick settings. $p(\mathcal{I}|\mathcal{S}_{n},\tau)$ can be seen as a rendering process studied in computer graphics. Here we are mainly interested in the silhouette, following the practice in textile in Cusick drape analysis [16]. Therefore, we replace $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)$ with a state transition function $\mathcal{S}_{i+1}=s(\mathcal{S}_{i},\tau)$ where $s$ is a deterministic function then $p(\mathcal{S}_{i+1}|\mathcal{S}_{i},\tau)=p(s(\mathcal{S}_{i},\tau)|\mathcal{S}_{i},\tau)=1$. Similarly, $p(\mathcal{I}|\mathcal{S}_{n},\tau)=p(r(\mathcal{S}_{n})|\mathcal{S}_{n},\tau)=1$ where $r$ is a rendering function. Therefore, Eq. 1 is transformed to:
@@ 4 Data Collection
The Cusick drape data is collected following the BS EN ISO 9073-9:2008 [1]. In every test, the warp/wale and weft/course directions are aligned across samples to ensure the same initial condition. In addition, we are able to reconstruct the 3D meshes of cloth drapes thanks to our patented drape meter. However, the 3D data are mainly for evaluation and the silhouette images are for learning. This is because not every Cusick drape meter can reconstruct 3D meshes out of cloth drapes. Being able to learn from the 2D silhouette images only is crucial in making our method applicable in real-world settings. In addition, we also measure the sample weight and calculate the average area density $\rho=(\sum_{i<13}^{i=1}m_{i})/(12\pi R^{2})$ where $m_{i}$ denotes the measured weight of cloth sample $i$ and $R=0.15m$. We also measure the thickness and include them in our dataset (refer to SM).
@@ 5 Experiments
Due to the nature of limited data, we conduct our experiments on small datasets. We use 5 representative types of cloths (each with 12 samples) that exhibit visually distinguishable drapability. For convenience, we name cloths as “material color”, e.g. Cotton Blue, Cotton Pink. To digitalize a cloth, we randomly select 1 out of 12 samples for training to learn their parameter distributions. During learning, we run 100 steps ($n=100$) for the forward simulation, with time step size $h=0.05s$ and the total time lapse is $5s$. To qualitatively evaluate the digitalization, we simulate garments made from different digitalized cloths and visually compare the simulated garments with their corresponding real cloths. To evaluate the model (i.e. drape fitting), we employ several metrics. For fitting capability, we use mean squared error (MSE) between the fitted and the ground truth drape images. We also use Hausdorff distance (H.Dis) between the simulated 3D mesh and the ground-truth (GT) mesh. Further, we use radius-angle graph (Fig. 6(a)[43] which is widely adopted for describing Cusick drape waves and comparing drape shapes. As generalization, we also test if our model learned on one sample can predict the physics of unseen cloth samples of the same type. https://youtu.be/ProN0y1bURY has more visual results.
@@ 5.2 Comparison
To our best knowledge, there is no similar method designed for exactly the same setting as ours. We therefore adopt [50] the closest methods as baselines because their method can also digitalize cloth by learning cloth physical parameters, albeit only taking simulation data as input. Since they only model homogeneous material, we refer to their model as HOMO. Further, we augment their model by making the cloth physical parameter learning element-wise, to enable learning heterogeneous materials. We call this model HETER for short. Both HOMO and HETER are deterministic models. So we use MSE as the loss function: $\mathcal{L}(\tau)=\sum_{i=1}^{L}\sum_{j=1}^{L}(\hat{\mathcal{I}}_{ij}-\mathcal{I}_{ij})^{2}$, for training. For comparison, we train HOMO and HETER with the same data (1 for training and the rest 11 for testing) as BDP. For BDP, we draw parameters from the learned parameter distributions for 1000 times and run Cusick drape simulations. Then, we select the best result to calculate the MSE and H. Dis.
@@
While our method is built on differentiable physics models for learning, i.e. derivative-based optimization, the traditional methods widely used in material science and physics are usually based on derivative-free optimization. So we also compare different learning strategies. Bayesian Optimization (BO) is a representative derivative-free optimizer which usually uses the Gaussian Process to approximate an unknown optimized objective function [12, 25]. However, the performance of vanilla BO drops drastically when the number of parameters is above 20 [24, 45]. There are over 200,000 parameters in the HETER cloth model. So we employ Random Embedding Bayesian Optimization (REMBO) [75] and Hashing-enhanced Subspace Bayesian Optimization (HeSBO) [57] as baselines. We use BoTorch’s [2] REMBO and HeSBO implementation and compare them with our BDP. Given the same drape silhouette, we run our method 1000 epochs (gain the best optimization result within 200 epochs), and the REMOB and HeSBO for 500 trials. Tab. 2 shows that, our derivative-based method is better with fewer optimization steps. More results can be found in the SM.